# AI-Assisted Rapid Crystal Structure Generation Towards a Target Local Environment

### v1

@@ III.4 Energy Ranking & Database
After the pre-relaxation, the valid structures need to be evaluated by the energy ranking. Given that many structures needs to be processed and it is computationally expensive to perform quantum mechanical calculations for each structure, we considered two alternatives, (i) a reactive force field (ReaxFF) Van Duin et al. (2001) via the GULP package Gale (1997), a bond-order-based force field that can efficiently model chemical reactions and interatomic interactions; and (ii) a pretrained MACE model Batatia et al. (2022) via ASE Larsen et al. (2017), which utilizes higher order equivariant message passing for fast and accurate predictions of interatomic interactions in a large chemical space. Both methods are significantly faster than conventional DFT calculations, making them suitable for large-scale screening of generated structures.
@@
Figure 4 compares energies calculated using MACE and ReaxFF models against reference DFT calculations performed with the VASP code Kresse and Furthmüller (1996) using the r2SCAN meta-GGA functional Furness et al. (2020a, b) and a KSPACING value of 0.15 Å-1. Generally, the MACE model shows a stronger correlation with DFT energies compared to ReaxFF, particularly for low-energy structures, making it slightly more suitable for our purposes. However, both models may stabilize structures into nonphysical configurations after pre-relaxation. Therefore, cross-checking generated structures with multiple models helps ensure physical validity. In our pipeline, we first use MACE for relaxation and energy ranking, then apply ReaxFF optimization to get the energy value as an alternative metric to infer the structural stability. According to our empirical observation, a feasible structure should be characterized by favorable energy ranking in both MACE and ReaxFF models.
@@ III.5 Approximated Time Costs
The rough timing information is shown in Table 2. Among all stages, the data collection and tabularization step can be done within several minutes on a single CPU. The time cost for model training is about 10-20 minutes using an H100 GPU for 250 epochs, and no more than 1 minute to generate 1k samples. The most significant time is spent on the pre-relaxation stage, as it needs to handle several hundred thousand samples with multiple iterations of optimization. On an AMD EPYC 9654P 96-Core Processor (each in 2.4 GHz), it takes 5-15 minutes per 1k samples and 8-16 hours for a total of 100k samples. Finally, the energy-based relaxation and ranking with MACE and ReaxFF require 5-20 minutes per 1k samples on the same 96-core processor, depending on the structure complexity and convergence behavior.
@@
Stage Device Speed
Data Augmentation Single CPU 3-5 mins
@@
Sampling H100 GPU 1~2 mins / 1k samples
Pre-relaxation AMD 96-core 5~15 mins / 1k samples
MACE-relaxation AMD 96-core 5~20 mins / 1k samples
ReaxFF-relaxation AMD 96-core 10~20 mins / 1k samples
@@ IV.1 Model Performances Evaluation
First, Table 3 shows statistics of unique sp2 structures from different generative models trained on the v1-60k dataset with 200~500 epochs. Compared to the original 140 unique structures in the SACADA dataset, all models except GAN-Cont. successfully regenerated over 100 training structures within 100k samples, indicating a strong capability to learn from the training data. If we allow generating more samples, we expect all the training structures would be reproduced by these models. More encouragingly, each model generated a substantial number of unique structures (GAN-Cont.: 2937, GAN-Dis.: 4514, VAE-Cont.: 4862, VAE-Dis.: 4289). The limited overlap between models, as found in Table 3, suggests that each model learn different statistical patterns from the training data, leading to diverse generated structures.
@@
Nevertheless, other VAE and GAN models demonstrate similar performances across various metrics. On average, we observe that around 98% success rate in generating valid crystal structures that satisfy the space group and Wyckoff site symmetry constraints, and 15-22% of these structures can be relaxed to adopt the desired sp2 bonding configuration. The number of unique structures ranges from 3,235 to 5,156 per 100k samples, indicating a good diversity in the generated crystal configurations.
For realistic application, the more important metric is the number of low-energy unique structures. The GAN-Dis-v2/v3 model achieved higher numbers of low-energy (cubic) structures as compared to the corresponding v1 model, suggesting that more explicit subgroup symmetries can promote the model’s learning efficacy. A similar trend is also observed for the VAE-continuous models. However, this is not the case for the VAE-Discrete models, in which the model trained on the v1-60k dataset produced the highest number of low-energy structures. This may be due to the fact that the VAE-Dis. model uses the noise sampled from continuous Gaussian distribution, but the generated results should be represented in discrete values. Therefore, the VAE models may have some intrinsic limitations in learning the discrete variables.
While it is overall difficult to make a definitive recommendation on a single best model for this task, we observe a general trend that most models can rapidly generate a large number of low-energy sp2 carbon allotropes with diverse structures, as compared to the original 140 training structures. This opens up new opportunities for effectively harnessing the power of AI generative models to explore the structural space of complicated materials. The improvements of model performance will be left for the future, as it is beyond the scope of this work.
@@ IV.2.1 Distribution of low-energy crystals by complexity
Figure 6: The screened low-dimensional sp2 carbon allotropes.. (a) Distribution of the energy of the generated structures grouped by dimensionality. (b-d) show three representative 0D, 1D, and 2D structures labeled by the Pearson symbol and the relative DFT-r2SCAN energies as compared to the ground state graphite structure.
@@ IV.2.2 The 0-2D low-energy sp2 carbon allotropes
Beyond the stacking graphite-type layers, we also identified several other planar configurations in different types of ring topology (e.g., mixed five-, seven, eight-membered rings, see Fig. S4 in the supplementary materials) with an energy spanning from 0 to 0.5 eV range. Additionally, several unique 1D carbon nanotube-like structures with varied diameters (e.g., a representative model 1D-hR108 in Figure 6c) have been observed in our results. Finally, we also found two large 0D allotropes (see 0D-cP176 in Figure 6d) featured by the combination of five, six and eight-membered rings that are similar to the well known C60 and C70 Haymet (1985). Notably, these 0D allotropes are calculated to possess lower energies (0.2-0.3 eV/atom) than C60 and C70. Hence, they have better stability if they can be synthesized in the experiment. These examples demonstrate our framework’s capability to explore diverse structural motifs across multiple dimensionalities.
@@ V Conclusions
In this work, we present LEGO-xtal, a novel framework that combines subgroup symmetry augmentation, descriptor-guided pre-relaxation, and advanced generative models to efficiently explore crystal structure space from limited training data. Starting from just 25 (140) known sp2 carbon structures, our approach successfully generated over 1,700 (21,786) diverse low-energy allotropes, including high-symmetry structures with large unit cells (up to 960 atoms) and unique 0D, 1D, and 2D configurations. This demonstrates the framework’s capability to discover complex crystal structures while maintaining physically realistic local environments.
@@
The success of our framework stems from the synergistic combination of subgroup symmetry augmentation and descriptor-guided pre-relaxation. While symmetry augmentation enables efficient exploration of the structural space through reduced discrete variables, pre-relaxation ensures the generated structures maintain desired local environments at minimal computational cost. Together, these strategies dramatically reduce the dimensionality of the crystal structure search space.
@@
Looking ahead, we aim to extend subgroup symmetry augmentation beyond single components and reference environments, enabling the generation of more complex structures with multiple components and varied local environments in a single model. This extension requires adapting pre-relaxation strategies to handle multiple reference environments, either by modifying the environmental loss function within traditional optimization routines Pickard (2025) or by employing more flexible diffusion models that incorporate local environment constraints during the training stage. These improvements will enable the generation of increasingly complex structures with higher dimensionality and more sophisticated local environments, which will be the focus of our future work.
@@ Acknowledgments
This research was sponsored by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, and the Established Program to Stimulate Competitive Research (EPSCoR) under the DOE Early Career Award No. DE-SC0024866, the UNC Charlotte’s seed grant for data science, as well as European Union (ERDF), Région Nouvelle Aquitaine, Poitiers Univeristy, and French government programs “Investissements d’Avenir" (EUR INTREE, reference ANR-18-EURE-0010) and PRC ANR MagDesign and TcPredictor. The computing resources are provided by ACCESS (TG-DMR180040) and High-Performance Computing Centre Adastra/CINES of GENCI (projects A0140807539 and A0160815101). During the preparation of this work, the authors used GitHub Copilot in order to improve the code readability and documentation. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ III.4 Energy Ranking & Database
After the pre-relaxation, the valid structures need to be evaluated by the energy ranking. Given that many structures needs to be processed and it is computationally expensive to perform quantum mechanical calculations for each structure, we considered two alternatives, (i) a reactive force field (ReaxFF) Van Duin et al. (2001) via the GULP package Gale (1997), a bond-order-based force field that can efficiently model chemical reactions and interatomic interactions; and (ii) a pretrained MACE model Batatia et al. (2022) via ASE Larsen et al. (2017), which utilizes higher order equivariant message passing for fast and accurate predictions of interatomic interactions in a large chemical space. Both methods are significantly faster than conventional DFT calculations, making them suitable for large-scale screening of generated structures. Although most of the pre-relaxed structures should be very close to the ground state, each energy model may describe their ground state slightly differently in terms of bond lengths and angles, as well as lattice constants. Therefore, it is still necessary to run relaxation, instead of single point energy evaluation, to obtain the ground state and accurate energy ranking.
@@
Figure 4 compares energies calculated using MACE and ReaxFF models against reference DFT calculations performed with the VASP code Kresse and Furthmüller (1996) using the r2SCAN meta-GGA functional Furness et al. (2020a, b) and a KSPACING value of 0.15 Å-1. In terms of the the root mean square error (RMSE) values, the MACE model (0.132 eV/atom) shows a better agreement with DFT energies compared to ReaxFF (0.147 eV/atom). Particularly, MACE describes better for the low-energy structures, making it slightly more suitable for our purposes. However, both models may stabilize structures into nonphysical configurations after pre-relaxation. Therefore, cross-checking generated structures with multiple models helps ensure physical validity. In our pipeline, we first use ReaxFF for relaxation and energy ranking, then apply MACE optimization to get the energy value as an alternative metric to infer the structural stability. According to our empirical observation, a feasible structure should be characterized by favorable energy ranking in both MACE and ReaxFF models.
@@ III.5 Approximated Time Costs
Table 2 summarizes the approximate timing for each stage in the LEGO-xtal workflow. All timings are based on processing 1,000 samples; for larger datasets, times scale linearly.
First, Data augmentation is performed on a single CPU and typically takes 0.2–0.5 minutes per 1,000 samples. Model training on an H100 GPU requires 10–20 minutes for 250 epochs, while sampling (generation of new structures) takes 1–2 minutes per 1,000 samples.
Next, Pre-relaxation using the SO(3) descriptor on an AMD EPYC 96-core processor (2.4 GHz) takes approximately 5.5 minutes per 1,000 samples. When SO(3) pre-relaxation is followed by GULP-ReaxFF or MACE, the combined times are 7.5 and 9.5 minutes per 1,000 samples, respectively. For comparison, running GULP-ReaxFF or MACE without pre-relaxation requires 20 and 120 minutes per 1,000 samples, respectively. This suggests that pre-relaxation can effectively reduce the time cost as compared to the energy-based optimization. As shown in the Supplementary Materials Section III-C, we also provide direct evidence to demonstrate that pre-relaxation also improve the success rate of finding the unique valid structures.
In addition, we have implemented Batchwise SO(3) pre-relaxation on an H100 GPU which is even faster, taking about 1.5 minute per 1,000 samples, although this approach is still under development and requires further optimization.
@@
Stage Device Speed (per 1k samples)
Data Augmentation Single CPU 0.2-0.5 mins
@@
Sampling H100 GPU 1~2 mins
Pre-relaxation AMD 96-core ~ 5.5 mins
Pre-relax + ReaxFF AMD 96-core ~ 5.5+1.5 mins
Pre-relax + MACE AMD 96-core ~ 5.5+2.5 mins
Pre-relaxation Batchwise H100 GPU ~1.5 mins
ReaxFF w/o Pre-relax AMD 96-core ~20 mins
MACE w/o Pre-relax AMD 96-core ~120 mins
@@ IV.1 Model Performances Evaluation
Table 3 shows statistics of unique sp2 structures from different generative models trained on the v1-60k dataset with 200~500 epochs. Compared to the original 140 unique structures in the SACADA dataset, all models except GAN-Cont. successfully regenerated over 100 training structures within 100k samples, indicating a strong capability to learn from the training data. If we allow generating more samples, we expect all the training structures would be reproduced by these models. More encouragingly, each model generated a substantial number of unique structures (GAN-Cont.: 2937, GAN-Dis.: 4514, VAE-Cont.: 4862, VAE-Dis.: 4289). The limited overlap between models, as found in Table 3, suggests that each model learn different statistical patterns from the training data, leading to diverse generated structures.
@@
Nevertheless, other VAE and GAN models demonstrate similar performances across various metrics. On average, we observe that around 98% success rate in generating valid crystal structures that satisfy the space group and Wyckoff site symmetry constraints, and 15-22% of these structures can be relaxed to adopt the desired sp2 bonding configuration. The number of unique structures ranges from 3,235 to 5,156 per 100k samples, indicating a good diversity in the generated crystal configurations. As shown in Fig. S4 of the Supplementary Materials, one can continue to try more samples to get new unique structures. In addition, our analysis in Section V-A of Supplementary Materials suggests that these metrics are clearly better than the pure random sampling approaches, suggesting the efficacy of the statistical model learning.
For realistic application, the more important metric is the number of low-energy unique structures. The GAN-Dis-v2/v3 model achieved higher numbers of low-energy (cubic) structures as compared to the corresponding v1 model, suggesting that more explicit subgroup symmetries can promote the model’s learning efficacy. A similar trend is also observed for the VAE-continuous models. However, this is not the case for the VAE-Discrete models, in which the model trained on the v1-60k dataset produced the highest number of low-energy structures. This may be due to the fact that the VAE-Dis. model uses the noise sampled from continuous Gaussian distribution, which may hinder the decoder from learning discrete, one-hot-like representations in the latent space, thus reducing generation fidelity on categorical outputs. A natural remedy is to adopt the Gumbel-Softmax reparameterization trick, which introduces a temperature-controlled, differentiable approximation to categorical sampling: starting from a high temperature to ensure smooth gradients, and annealing towards a low but non-zero $\tau$ to encourage near-discrete latent codes while preserving differentiability Jang, Gu, and Poole (2017). Additionally, balancing the KL-divergence and reconstruction terms via $\beta$-VAE Higgins et al. (2017) or KL-annealing Fu et al. (2019) strategies can prevent posterior collapse and improve latent expressiveness. Finally, in scenarios with severely imbalanced categorical outputs, applying class-weighted Cross-Entropy or Focal Loss Lin et al. (2017) can stabilize training by emphasizing hard or rare classes.
Unlike VAE models that explicitly reconstruct inputs, GANs rely solely on adversarial feedback. This enables the discriminator to learn the joint distribution of the representations without being constrained to memorizing input patterns, thereby enhancing the generator’s ability to generalize. This effect is particularly evident in Table 4, where the GAN achieves stronger performance on the discrete dis-dataset compared to the Cont dataset. The smaller and more structured joint distribution space of discrete features makes it easier for the GAN to capture meaningful generalizable patterns, rather than overfitting to training samples.
While it is difficult to make a definitive recommendation on a single best model for this task, we observe a general trend that most models can rapidly generate a large number of low-energy sp2 carbon allotropes with diverse structures, as compared to the original 140 training structures. Importantly, Table 4 suggests that the more data is included, the chance of generating unique structures ($N_{\text{unique}}$) becomes higher regardless choice of generative models. This is likely to be associated with the increment of unique symmetries (see extended discussion in Section V-B of the Supplementary Materials). Overall, the results highlight new opportunities for effectively harnessing the power of AI generative models to explore the structural space of complicated materials. The improvements of model performance will be left for the future, as it is beyond the scope of this work.
@@ IV.2.1 Distribution of low-energy crystals by complexity
Figure 6: The screened low-dimensional sp2 carbon allotropes. (a) Distribution of the energy of the generated structures grouped by dimensionality. (b-d) show three representative 0D, 1D, and 2D structures labeled by the Pearson symbol and the relative DFT-r2SCAN energies as compared to the ground state graphite structure.
@@ IV.2.2 The 0-2D low-energy sp2 carbon allotropes
Beyond the stacking graphite-type layers, we also identified several other planar configurations in different types of ring topology (e.g., mixed five-, seven, eight-membered rings, see Section VI and Fig. S6 in the Supplementary Materials) with an energy spanning from 0 to 0.5 eV range. Additionally, several unique 1D carbon nanotube-like structures with varied diameters (e.g., a representative model 1D-hR108 in Figure 6c) have been observed in our results. Finally, we also found two large 0D allotropes (see 0D-cP176 in Figure 6d) featured by the combination of five, six and eight-membered rings that are similar to the well known C60 and C70 Haymet (1985). Notably, these 0D allotropes are calculated to possess lower energies (0.2-0.3 eV/atom) than C60 and C70. Hence, they have better stability if they can be synthesized in the experiment. These examples demonstrate our framework’s capability to explore diverse structural motifs across multiple dimensionalities.
@@ V Conclusions
In this work, we present LEGO-xtal, a novel framework that combines subgroup symmetry augmentation, descriptor-guided pre-relaxation, and modern generative components (GAN and VAE) to efficiently explore crystal structure space from limited training data. Starting from just 25 (140) known sp2 carbon structures, our approach successfully generated over 1,700 (21,786) diverse low-energy allotropes, including high-symmetry structures with large unit cells (up to 960 atoms) and unique 0D, 1D, and 2D configurations. This demonstrates the framework’s capability to discover complex crystal structures from a small dataset while maintaining physically desirable local environments. This clearly distinguishes our approach from previous methods Zhu et al. (2024); Merchant et al. (2023); Xie et al. (2021); Jiao et al. (2023); Zeni et al. (2025); Levy et al. (2024); Jiao et al. (2024); Miller et al. (2024); Zhong et al. (2025); Antunes, Butler, and Grau-Crespo (2024); Gruver et al. (2024); Cao et al. (2024); Kazeev et al. (2025) that focus on generation of relatively small crystal systems from a large dataset without the explicit constraints on the local chemical environment.
@@
The success of our framework stems from the synergistic combination of subgroup symmetry augmentation and descriptor-guided pre-relaxation. While symmetry augmentation enables efficient exploration of the structural space through reduced discrete variables, pre-relaxation ensures the generated structures maintain desired local environments at minimal computational cost. Together, these strategies dramatically reduce the dimensionality of the crystal structure search space. Importantly, the subgroup-based data augmentation strategy introduced in our work is broadly applicable. In the supplementary materials Section IV-C, we have provided a universal interface to systemically search for subgroup symmetries, which could be integrated into existing symmetry-aware generative models to enhance their generalization capabilities Zhu et al. (2024); Levy et al. (2024); Jiao et al. (2024); Antunes, Butler, and Grau-Crespo (2024); Cao et al. (2024); Kazeev et al. (2025).
Regarding the choice of generative models, this study explored both VAE and GAN architectures that offer a balance between simplicity and effectiveness. Despite their relative simplicity, both remain core building blocks in today’s state-of-the-art pipelines (e.g., Stable Diffusion Rombach et al. (2022), SV4D 2.0 Yao et al. (2025), MaskGIT Chang et al. (2022)). We hope that LEGO-xtal will serve as a solid starting point, from which we can build toward integrating more advanced generative modeling techniques more advanced paradigms such as latent diffusion, masked modeling, or auto-regressive architectures in future work.
Due to the use of symmetrized tabular representation, this approach is likely to intrinsically favor high symmetry structures. For instance, our current models do not consider the generation of more than 8 atoms in the unit cell with PP1 symmetry. While the assumption of no low-energy complex structures in low symmetry is true for the case of sp2 carbon allotropes, it may not be suitable to describe other systems where low-symmetry crystals appear more often. In that case, we recommend the use of the generative models without symmetry constraints Zeni et al. (2025); Jiao et al. (2023); Miller et al. (2024).
@@
Finally, we aim to extend subgroup symmetry augmentation beyond single components and reference environments, enabling the generation of more complex structures with multiple components and varied local environments in a single model. This extension requires (1) adapting pre-relaxation strategies to handle multiple reference environments, either by modifying the environmental loss function within traditional optimization routines Pickard (2025) or by employing more flexible diffusion models that incorporate local environment constraints during the training stage and (2) the support of generative models on Wyckoff sites labeled with different elements or local environments. Conceptually, the use of descriptor-guided pre-relaxation and subgroup augmentation is expected to be even more advantageous to handle more complex compositional systems. These improvements will enable the generation of increasingly complex structures with higher dimensionality and more sophisticated local environments, which will be the focus of our future work.
@@ Acknowledgments
This research was sponsored by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, and the Established Program to Stimulate Competitive Research (EPSCoR) under the DOE Early Career Award No. DE-SC0024866, the UNC Charlotte’s seed grant for data science, as well as European Union (ERDF), Région Nouvelle Aquitaine, Poitiers Univeristy, and French government programs “Investissements d’Avenir" (EUR INTREE, reference ANR-18-EURE-0010) and PRC ANR MagDesign and TcPredictor. The computing resources are provided by ACCESS (TG-DMR180040) and High-Performance Computing Centre Adastra/CINES of GENCI (projects A0140807539 and A0160815101). We also acknowledge the reviewers for their insightful comments and suggestions during the review stage. During the preparation of this work, the authors used GitHub Copilot in order to improve the code readability and documentation. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.